I have a work sheet with quadratic equations and i don't understand it at all. It read's: Solve the equation by factoring, 3xsquared+9x-12=0.(how do i factor this where do i start?)my teacher showed me how but i can't remember can you please help me?
Answer by mukhopadhyay(490)

3x^2+9x-12=0
=>3x^2+12x-3x-12=0
=>3x(x+4)-3(x+4)=0
=>(x+4)(3x-3)=0
=>(x+4)=0 or (3x-3)=0
=>x=-4 or 3x=3
=>x=-4 or x=1
Answer: x=-4 or x=1